Find the vertices and foci of the ellipse given by the equation 4x^2 + 9x^2 =144
Answer by ichigo449(30)

Put into canonical form: x^2/36 + y^2/16 = 1. Now the vertices are (0,6) and (0,-6). The distance to the focus is: sqrt(36-16) = sqrt(20). The foci of the ellipse are then: (sqrt(20),0) and (-sqrt(20),0). Have a nice day.